# Parameter Invariance Analysis of Moment Equations Using Dulmage-Mendelsohn Decomposition

### v1

@@ IV Application example
Consider the chemical reaction network in Fig. 2(a), where the network consists of $n=6$ molecular species and $m=11$ reactions. Here, we aim to identify the invariant stationary moments under the variation of $\theta_{3}$, i.e., $\mathcal{M}_{3}$ in eq. (3).
For this purpose, we first formulate the stationary moment equation (5) up to the second order moments, which results in 27 equations with 32 variables. Next, we apply the DM decomposition to the structure matrix $\tilde{A}$ to extract the square matrix $\tilde{A}_{d}$ (Lemma 1). The resulting matrix is shown in Fig. 2(b), which indeed becomes the upper-triangular form. In particular, the square matrix $\tilde{A}_{d}$ with $N_{d}=7$ dimensions with $\eta=7$ diagonal blocks (scalars) is successfully extracted from the underdetermined moment equation. Applying the same row and column permutations, i.e., $P$ and $Q$ in Lemma 1, to eq. (8), we further obtain $\Psi\bm{\nu}_{\Psi}=\bm{b}_{\Psi}$ in eq. (12) (Lemma 2 and Appendix -A). Figure 2(c) shows the welldetermined part of the equation $\Psi\bm{\nu}_{\Psi}=\bm{b}_{\Psi}$, which becomes $142N_{d}=$ 14 dimensions in the upper-triangular form. Finally, using Theorem 1, we explore the set of moments $\mathcal{M}_{3}$. Specifically, a graph $G(\mathcal{V},\mathcal{E})$ with 14 nodes with two source nodes at $V_{8}$ and $V_{14}$ is defined since $b_{8}\neq 0$ and $b_{14}\neq 0$. Then, the reachable states are examined by tracing all reachable nodes from the sources (see the orange lines in Fig. 2(c)).
Figure 2: Application example. (a) a network of feedback reactions, (b) DM decomposition of $\tilde{A}$, and (c) identification of invariant moments based on the backward substitution of eq. (12)
From these results, it is concluded that $\mathbb{E}[x_{1}x_{6}],\mathbb{E}[x_{5}],\mathbb{E}[x_{4}],$ and $\mathbb{E}[x_{3}]$ are invariant moments $\mathcal{M}_{3}$ with respect to the variation of $\theta_{3}$. These results demonstrate the systematic and computationally efficient feature of the proposed approach, which leverages the DM decomposition of the coefficient matrix along with the unreachability analysis of the graph. The identified invariant moments provide deeper insights into homeostatic behaviors of biomolecular systems.
@@ V Conclusion
In this paper, we have proposed a systematic method for analyzing the parameter invariance of stationary moments in general chemical reaction networks by leveraging the non-zero structure of the moment equations. First, by applying the Dulmage-Mendelsohn (DM) decomposition to the moment equations, we have extracted welldetermined subequations from the underdetermined equations and have revealed the hierarchical structure among stationary moments. Furthermore, by performing unreachability analysis on a directed graph based on this structure, we have demonstrated that stationary moments whose partial derivatives with respect to parameters are structurally zero can be systematically identified. The proposed method enables rigorous parameter invariance analysis for a broad class of stochastic chemical systems, offering new insights into the homeostasis of chemical reaction networks.

### v2

@@ IV Application example
Consider the chemical reaction network shown in Fig. 2(a), which has an Antithetic Integral Feedback (AIF) structure formed by a sequestration reaction[16, 4], and consists of $n=6$ molecular species and $m=12$ reactions. Here, we aim to identify the invariant stationary moments under the variation of $\theta_{4}$, i.e., $\mathcal{M}_{4}$. Formulating the stationary moment equations up to $\sum^{n}_{i=1}\alpha_{i}=1$, we have
$\!\!\!\!\begin{bmatrix}0&\!\!0&\!\!\!0&\!\!\!0&\!\!\!0&\!\!\!\!-\theta_{11}&\!\!\!\!0\\ \theta_{2}&\!\!\theta_{3}&\!\!\!0&\!\!\!0&\!\!\!\theta_{9}&\!\!\!\!0&\!\!\!\!-\theta_{3}\\ 0&\!\!\theta_{4}&\!\!\!-\theta_{5}&\!\!\!\theta_{6}&\!\!\!0&\!\!\!\!0&\!\!\!\!0\\ 0&\!\!0&\!\!\!\theta_{5}&\!\!\!-\theta_{6}&\!\!\!0&\!\!\!\!0&\!\!\!\!0\\ 0&\!\!0&\!\!\!\theta_{7}&\!\!\!0&\!\!\!-\theta_{8}&\!\!\!\!0&\!\!\!\!0\\ 0&\!\!0&\!\!\!0&\!\!\!0&\!\!\!\theta_{10}&\!\!\!\!-\theta_{11}&\!\!\!\!0\\ \end{bmatrix}\!\!\begin{bmatrix}\!\mathbb{E}[x_{1}]\!\\ \!\mathbb{E}[x_{2}]\!\\ \!\mathbb{E}[x_{3}]\!\\ \!\mathbb{E}[x_{4}]\!\\ \!\mathbb{E}[x_{5}]\!\\ \!\mathbb{E}[x_{1}x_{6}]\!\\ \!\mathbb{E}[x^{2}_{2}]\!\end{bmatrix}\!\!=\!\!\begin{bmatrix}\!-\theta_{1}\!\\ \!0\!\\ \!0\!\\ \!-\theta_{12}\!\\ \!0\!\\ \!0\!\end{bmatrix}\!\!\!$ (15)
In this form, the system is underdetermined with six equations and seven variables. Applying Lemma 2 and extracting the welldetermined subequations, we obtain eq. (12), of which the sparsity pattern is shown Fig. 2(b). Finally, using Theorem 1, we explore invariant moments $\mathcal{M}_{4}$. Specifically, a graph $\mathcal{G}(\mathcal{V},\mathcal{E})$ with 10 nodes with two source nodes at $V_{7}$ and $V_{10}$ is defined since $b_{7}\neq 0$ and $b_{10}\neq 0$. Then, the reachable states are examined by tracing all reachable nodes from the sources (see the orange lines in Fig. 2(b)).
Figure 2: Application example. (a) a network of feedback reactions, (b) identification of invariant moments based on the backward substitution of eq. (12).
The result shows that $\mathbb{E}[x_{1}x_{6}],\mathbb{E}[x_{5}],\mathbb{E}[x_{3}],$ and $\mathbb{E}[x_{4}]$ are invariant moments $\mathcal{M}_{4}$ with respect to the variation of $\theta_{4}$.
These results agree with the existing results for the AIF systems [16, 17], which guarantee the invariance of $\mathbb{E}[x_{5}]$ and $\mathbb{E}[x_{1}x_{6}]$, while the proposed method enables to further discover the invariance of $\mathbb{E}[x_{3}]$ and $\mathbb{E}[x_{4}]$. This demonstrates that the proposed method can systematically identify invariant moments with respect to arbitrary parameters using a polynomial-time algorithm, addressing a problem that is generally considered challenging due to the infinite dimensionality of the moment equations. The identified invariant moments provide deeper insights into the homeostatic behavior of biomolecular systems, advancing both the fundamental understanding of biological robustness and the engineering of robust gene circuits for synthetic biology applications.
@@ V Conclusion
In this paper, we have proposed a systematic method for analyzing the parameter invariance of stationary moments in general chemical reaction networks by leveraging the non-zero structure of the moment equations. Specifically, we revealed the hierarchical structure among stationary moments using the Dulmage-Mendelsohn (DM) decomposition, and demonstrated that invariant stationary moments can be systematically identified by performing unreachability analysis on a graph based on this structure. The proposed method provides a computationally efficient framework that offers new insights into the homeostatic behavior in biomolecular systems. Future work will focus on increasing the number of identifiable invariant moments by incorporating parameter dependencies, as well as extending the framework toward the design and control problems of target moments.